# references.bib

### v1

@@ 4 Conclusion
Many avenues for future research remain open. First, our results were limited to small and relatively simple hypergraph structures, with most results computed for hypergraphs on $n=6$ nodes up to a maximum of $n=9$, and the largest number of hyperedges $m=20$. Relatedly, all of the node co-occurrence projection matrices we studied were sampled based on a $k$-uniform hypergraph model with a homogeneous node hyperdegree distribution. This leaves open the possibility of studying larger hypergraph structures as well as more complex synthetic models that explicitly incorporate features like hyperdegree and hyperedge size heterogeneity, node clustering, hyperedge overlap, and modular or community structure.
@@
It may also be fruitful to further explore how properties of node co-occurrence matrices can be used to directly predict properties of their sets of twins. Here we have taken some first steps by evaluating the size of the search tree; number of twins, twin isomorphism classes, and pairs of Gram Mates; diversity of hyperedge size distributions; and pairwise distances within sets of twins. However, we did not identify a single measure that can be computed from a node co-occurrence matrix and predict the complexity of the set of twins. Such a measure could in turn be used to immediately evaluate, at least heuristically, whether the node co-occurrence matrix of a given hypergraph may be more or less susceptible to problems associated with non-uniqueness. It is clear from our results that properties such as the density of cliques in the projection matrix and the worst case search tree size are important factors, but neither is strictly predictive.
We have also only analyzed pairwise hypergraph distances using a small selection of potential measures, which could be expanded in the future [vasilyeva_distances_2023, qin_explainable_2023]. It may also be promising to explore the possibility of using sets of twins to design a within-ensemble hypergraph distance [hartle_network_2020].
Finally, future work may focus specifically on the relationship between node co-occurrence matrices, sets of twins, and dynamical outcomes. Previous work has shown that the existence of twins matters [salova_analyzing_2021, ferraz_de_arruda_phase_2021]. The ability to enumerate these twin hypergraphs opens the door to a systematic approach to understanding the impact of non-uniqueness on dynamics at both lower and higher orders.
@@ Acknowledgments
The authors acknowledge support from the EPSRC Grant EP/V03474X/1. TL acknowledges Karel Devriendt, Brennan Klein, Maxime Lucas, Nicola Pedreschi, Rohit Sahasrabuddhe and Jason P. Smith for helpful conversations about the ideas in this paper. TL also acknowledges open source code made available by the developers of many projects including NumPy [harris2020array], SciPy [2020SciPy-NMeth], NetworkX [SciPyProceedings_11], MatPlotLib [Hunter_2007], and compleX Group Interactions (XGI) [Landry_XGI_A_Python_2023]. TL acknowledges support from a grant from the Fund for Energy Research with Corporate Partners administered by the Andlinger Center for Energy and the Environment at Princeton University as well as the School of Engineering and Applied Science (SEAS) Seed Grant at Princeton University.

### v2

@@ 4 Conclusion
Many avenues for future research remain open. First, our results were limited to small and relatively simple hypergraph structures, with most results computed for hypergraphs on $n=6$ nodes up to a maximum of $n=9$, and the largest number of hyperedges $m=20$. Relatedly, all of the node co-occurrence projection matrices we studied were sampled based on a $k$-uniform hypergraph model with a homogeneous node hyperdegree distribution. This leaves open the possibility of studying larger hypergraph structures as well as more complex random models that explicitly incorporate features like hyperdegree and hyperedge size heterogeneity, node clustering, hyperedge overlap, and modular or community structure.
@@
It may also be fruitful to further explore how properties of node co-occurrence matrices can be used to directly predict properties of their sets of twins. Here we have taken some first steps by evaluating the size of the search tree; the clique approximation; the number of twins, twin isomorphism classes, and pairs of Gram Mates; the diversity of hyperedge size distributions; and pairwise distances within sets of twins. However, we did not identify a single measure that can be computed from a node co-occurrence matrix and predict the complexity of the set of twins, nor did we find a single comparative measure that can correctly rank matrices by their twin set size without enumerating it. Such a measure could be used in conjunction with a null model to immediately evaluate, at least heuristically, whether the node co-occurrence matrix of a given hypergraph may be more or less susceptible to problems associated with non-uniqueness. It is clear from our results that properties such as the density of cliques in the projection matrix, the worst case search tree size, and the clique and hyperedge size distributions are all important factors, but we did not find any one of these factors to be strictly predictive.
We have also only analyzed hypergraph distances among sets of twins using a small selection of potential measures, which could be expanded in the future [vasilyeva_distances_2023, qin_explainable_2023]. It may also be promising to explore the possibility of using sets of twins to design a within-ensemble hypergraph distance, similar to the distance defined by Hartle et al. [hartle_network_2020].
Finally, future work may focus specifically on the relationship between node co-occurrence matrices, sets of twins, and dynamical outcomes. Previous work has shown that the existence of twins matters [salova_analyzing_2021, ferraz_de_arruda_phase_2021]. The ability to enumerate these twin hypergraphs opens the door for systematic approaches to understanding the impact of non-uniqueness on dynamics at both lower and higher orders.
@@ Acknowledgments
The authors acknowledge support from the EPSRC Grant EP/V03474X/1. TL acknowledges Karel Devriendt, Brennan Klein, Maxime Lucas, Jelle Oostveen, Nicola Pedreschi, Rohit Sahasrabuddhe and Jason P. Smith for helpful conversations about the ideas in this paper. TL also acknowledges open source code made available by the developers of many projects including NumPy [harris2020array], SciPy [2020SciPy-NMeth], NetworkX [SciPyProceedings_11], MatPlotLib [Hunter_2007], and compleX Group Interactions (XGI) [Landry_XGI_A_Python_2023]. TL acknowledges support from a grant from the Fund for Energy Research with Corporate Partners administered by the Andlinger Center for Energy and the Environment at Princeton University as well as the School of Engineering and Applied Science (SEAS) Seed Grant at Princeton University.
@@
Appendix A Rank Correlation Coefficients
In Table 1, we show for each value of $m$ the number of unique up to isomorphism node co-occurrence matrices, as well as the numerical values of Kendall’s $\tau$ computed for the clique and worst case tree size approximations and the associated $p$-values corresponding to the data shown in the rightmost plot of Figure 6.
Clique Approximation Worst Case Tree Size
$m$ matrices $\tau$ $p$-value $\tau$ $p$-value
3 3 1.000000 0.157299 1.000000 0.157299
4 15 0.878833 0.000012 0.717409 0.000511
5 37 0.600001 0.000000 0.535934 0.000005
6 88 0.533118 0.000000 0.494202 0.000000
7 155 0.427996 0.000000 0.383663 0.000000
8 243 0.403728 0.000000 0.331698 0.000000
9 303 0.432917 0.000000 0.312208 0.000000
10 342 0.461773 0.000000 0.322247 0.000000
11 304 0.535480 0.000000 0.335398 0.000000
12 245 0.605213 0.000000 0.347856 0.000000
13 159 0.663137 0.000000 0.363363 0.000000
14 94 0.696488 0.000000 0.366120 0.000000
15 43 0.740864 0.000000 0.371827 0.000761
16 21 0.647619 0.000011 0.433594 0.008701
17 7 0.809524 0.010714 0.514344 0.116852
18 3 1.000000 0.333333 0.000000 1.000000
Table 1: Kendall’s Tau rank correlation coefficients and p-values for points from Figure 6.